OpenStudy (anonymous):

The first term of a finite arithmetic sequence is 20 and the last is 50. The sum of this series is 385. How many terms does it have? (please show working)

OpenStudy (mertsj):

Sum = n(1st term + last term)/2 385 = n(20+50)/2 385 = 35n n=11

OpenStudy (anonymous):

Thanks alot

OpenStudy (mertsj):

yw